Gain (V/V) R Setting Totals Averages Sample 1 Sample 2 Sample 3 4 ап 7.8 8.1 7.9 3 5.2 6.0 4.3 = 359.3 i=1 j=1 2 4.4 6.9 3.8Use Fishers Least Significant Difference to determine which, if any, pairs of resistance settings significantly affect gainFormulae: an SST = i=1 j=1 -ΣΣνή with an - 1 degrees of freedom SStr = Ś ył_y? with a - 1 degrees of freedom n N i=1 SSE = SSf0.25.01.12 3 4 5 6 7 1 2 3 3 4 1 5.83 2.57 2.02 1.81 1.69 1.62 152 1.57 154 1.54 151 1.SI 140 1.49 1.47 1.46 1.45 1.44 1.430 TABLE. V Percentage Points ta of the t Distribution O. .40 .25 .05 2 .325 .289 .277 .271 .267 1.000 .816 .765 .10 3.078 1.8
Answer #1

The following table is obtained:

Group 1 Group 2 Group 3
7.8 8.1 7.9
5.2 6.0 4.3
4.4 6.9 3.8
2.0 1.7 0.8
Sum = 19.4 22.7 16.8
Average = 4.85 5.675 4.2
\sum_i X_{ij}^2 =∑i​Xij2​= 111.24 152.11 95.98
St. Dev. = 2.391 2.786 2.911
SS = 17.15 23.2875 25.42
n = 4 4 4

The total sample size is N = 12N=12. Therefore, the total degrees of freedom are:

dftotal​=12−1=11

Also, the between-groups degrees of freedom are dfbetween​=3−1=2, and the within-groups degrees of freedom are:

dfwithin​=dftotal​−dfbetween​=11−2=9

First, we need to compute the total sum of values and the grand mean. The following is obtained

∑​Xij​=19.4+22.7+16.8=58.9

Also, the sum of squared values is

∑​Xij2​=111.24+152.11+95.98=359.33

Based on the above calculations, the total sum of squares is computed as follows

2 SSIotal = ΣΧ ΣΑ; (ΣΑ, = 359.33 58.92 12 - 70.229 1, 2.]

The within sum of squares is computed as shown in the calculation below:

SS_{within} = \sum SS_{within groups} = 17.15+23.2875+25.42 = 65.858

SS_{treatment} = SS_{total}-SS_{within} =70.229-65.858=4.372

MS_{between} = \frac{SS_{between}}{df_{between}} = \frac{ 4.372}{ 2} = 2.186

MS_{within} = \frac{SS_{within}}{df_{within}} = \frac{ 65.858}{ 9} = 7.318

Finally, with having already calculated the mean sum of squares, the F-statistic is computed as follows:

F = \frac{MS_{between}}{MS_{within}} = \frac{ 2.186}{ 7.318} = 0.299

Source of variation SS df MS F
treatment 4.372 2 2.186 0.299
error 65.858 9 7.318
total 70.229 11

Tukey HSD results

treatments
pair
Tukey HSD
Q statistic
Tukey HSD
p-value
Tukey HSD
inferfence
A vs B 0.6100 0.8999947 insignificant
A vs C 0.4806 0.8999947 insignificant
B vs C 1.0905 0.7184406 insignificant

No pair significantly effect resistance settings.

please rate my answer and comment for doubts.